Trinity Top Ten 2001

Round 1 Due Date : [July 27]

DURC

Mark Dennehy

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]

DURC

Mark Dennehy

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 2 Due Date : [August 10]

DURC

Mark Dennehy

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 2 Due Date : [August 10]

DURC

Mark Dennehy

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 3 Due Date : [August 24]

DURC

Mark Dennehy

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 3 Due Date : [August 24]

DURC

Mark Dennehy

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 4

Due Date : [September 7]

DURC

Mark Dennehy

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 4

Due Date : [September 7]

DURC

Mark Dennehy

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 5

Due Date : [September 21]

DURC

Mark Dennehy

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 5

Due Date : [September 21]

DURC

Mark Dennehy

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 6 Due Date : [October 5]

DURC

Mark Dennehy

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 6 Due Date : [October 5]

DURC

Mark Dennehy

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 7 Due Date : [October 19]

DURC

Mark Dennehy

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 7 Due Date : [October 19]

DURC

Mark Dennehy

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 8 Due Date : [November 2]

DURC

Mark Dennehy

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 8 Due Date : [November 2]

DURC

Mark Dennehy

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 9

Due Date : [November 16]

DURC

Mark Dennehy

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 9

Due Date : [November 16]

DURC

Mark Dennehy

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 10

Due Date : [November 30]

DURC

Mark Dennehy

25yd Prone Rifle Trinity Top Ten 2001

Round 10

Due Date : [November 30]

DURC

Mark Dennehy

25yd Prone Rifle

Place behind first bull of target

Round 1 Due Date : [July 27]

DURC

Alan Smyth

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]

DURC

Alan Smyth

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 2 Due Date : [August 10]

DURC

Alan Smyth

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 2 Due Date : [August 10]

DURC

Alan Smyth

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 3 Due Date : [August 24]

DURC

Alan Smyth

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 3 Due Date : [August 24]

DURC

Alan Smyth

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 4

Due Date : [September 7]

DURC

Alan Smyth

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 4

Due Date : [September 7]

DURC

Alan Smyth

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 5

Due Date : [September 21]

DURC

Alan Smyth

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 5

Due Date : [September 21]

DURC

Alan Smyth

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 6 Due Date : [October 5]

DURC

Alan Smyth

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 6 Due Date : [October 5]

DURC

Alan Smyth

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 7 Due Date : [October 19]

DURC

Alan Smyth

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 7 Due Date : [October 19]

DURC

Alan Smyth

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 8 Due Date : [November 2]

DURC

Alan Smyth

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 8 Due Date : [November 2]

DURC

Alan Smyth

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 9

Due Date : [November 16]

DURC

Alan Smyth

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 9

Due Date : [November 16]

DURC

Alan Smyth

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 10

Due Date : [November 30]

DURC

Alan Smyth

25yd Prone Rifle Trinity Top Ten 2001

Round 10

Due Date : [November 30]

DURC

Alan Smyth

25yd Prone Rifle

Place behind first bull of target

Round 1 Due Date : [July 27]

DURC

John Keeney

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]

DURC

John Keeney

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 2 Due Date : [August 10]

DURC

John Keeney

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 2 Due Date : [August 10]

DURC

John Keeney

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 3 Due Date : [August 24]

DURC

John Keeney

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 3 Due Date : [August 24]

DURC

John Keeney

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 4

Due Date : [September 7]

DURC

John Keeney

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 4

Due Date : [September 7]

DURC

John Keeney

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 5

Due Date : [September 21]

DURC

John Keeney

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 5

Due Date : [September 21]

DURC

John Keeney

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 6 Due Date : [October 5]

DURC

John Keeney

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 6 Due Date : [October 5]

DURC

John Keeney

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 7 Due Date : [October 19]

DURC

John Keeney

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 7 Due Date : [October 19]

DURC

John Keeney

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 8 Due Date : [November 2]

DURC

John Keeney

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 8 Due Date : [November 2]

DURC

John Keeney

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 9

Due Date : [November 16]

DURC

John Keeney

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 9

Due Date : [November 16]

DURC

John Keeney

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 10

Due Date : [November 30]

DURC

John Keeney

25yd Prone Rifle Trinity Top Ten 2001

Round 10

Due Date : [November 30]

DURC

John Keeney

25yd Prone Rifle

Place behind first bull of target

Round 1 Due Date : [July 27]

DURC

Peter Byrne

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]

DURC

Peter Byrne

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 2 Due Date : [August 10]

DURC

Peter Byrne

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 2 Due Date : [August 10]

DURC

Peter Byrne

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 3 Due Date : [August 24]

DURC

Peter Byrne

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 3 Due Date : [August 24]

DURC

Peter Byrne

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 4

Due Date : [September 7]

DURC

Peter Byrne

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 4

Due Date : [September 7]

DURC

Peter Byrne

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 5

Due Date : [September 21]

DURC

Peter Byrne

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 5

Due Date : [September 21]

DURC

Peter Byrne

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 6 Due Date : [October 5]

DURC

Peter Byrne

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 6 Due Date : [October 5]

DURC

Peter Byrne

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 7 Due Date : [October 19]

DURC

Peter Byrne

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 7 Due Date : [October 19]

DURC

Peter Byrne

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 8 Due Date : [November 2]

DURC

Peter Byrne

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 8 Due Date : [November 2]

DURC

Peter Byrne

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 9

Due Date : [November 16]

DURC

Peter Byrne

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 9

Due Date : [November 16]

DURC

Peter Byrne

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 10

Due Date : [November 30]

DURC

Peter Byrne

25yd Prone Rifle Trinity Top Ten 2001

Round 10

Due Date : [November 30]

DURC

Peter Byrne

25yd Prone Rifle

Place behind first bull of target

Round 1 Due Date : [July 27]

DURC

Conor McDermottroe

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]

DURC

Conor McDermottroe

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 2 Due Date : [August 10]

DURC

Conor McDermottroe

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 2 Due Date : [August 10]

DURC

Conor McDermottroe

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 3 Due Date : [August 24]

DURC

Conor McDermottroe

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 3 Due Date : [August 24]

DURC

Conor McDermottroe

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 4

Due Date : [September 7] DURC

Conor McDermottroe

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 4

Due Date : [September 7] DURC

Conor McDermottroe

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 5

Due Date : [September 21] DURC

Conor McDermottroe

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 5

Due Date : [September 21] DURC

Conor McDermottroe

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 6 Due Date : [October 5]

DURC

Conor McDermottroe

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 6 Due Date : [October 5]

DURC

Conor McDermottroe

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 7 Due Date : [October 19]

DURC

Conor McDermottroe

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 7 Due Date : [October 19]

DURC

Conor McDermottroe

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 8 Due Date : [November 2]

DURC

Conor McDermottroe

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 8 Due Date : [November 2]

DURC

Conor McDermottroe

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 9

Due Date : [November 16] DURC

Conor McDermottroe

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 9

Due Date : [November 16] DURC

Conor McDermottroe

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 10

Due Date : [November 30] DURC

Conor McDermottroe

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 10

Due Date : [November 30] DURC

Conor McDermottroe

25yd Prone Rifle

Place behind first bull of target

Round 1 Due Date : [July 27]

DURC

Adrian Costigan

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]

DURC

Adrian Costigan

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 2 Due Date : [August 10]

DURC

Adrian Costigan

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 2 Due Date : [August 10]

DURC

Adrian Costigan

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 3 Due Date : [August 24]

DURC

Adrian Costigan

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 3 Due Date : [August 24]

DURC

Adrian Costigan

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 4

Due Date : [September 7]

DURC

Adrian Costigan

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 4

Due Date : [September 7]

DURC

Adrian Costigan

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 5

Due Date : [September 21]

DURC

Adrian Costigan

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 5

Due Date : [September 21]

DURC

Adrian Costigan

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 6 Due Date : [October 5]

DURC

Adrian Costigan

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 6 Due Date : [October 5]

DURC

Adrian Costigan

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 7 Due Date : [October 19]

DURC

Adrian Costigan

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 7 Due Date : [October 19]

DURC

Adrian Costigan

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 8 Due Date : [November 2]

DURC

Adrian Costigan

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 8 Due Date : [November 2]

DURC

Adrian Costigan

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 9

Due Date : [November 16]

DURC

Adrian Costigan

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 9

Due Date : [November 16]

DURC

Adrian Costigan

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 10

Due Date : [November 30]

DURC

Adrian Costigan

25yd Prone Rifle Trinity Top Ten 2001

Round 10

Due Date : [November 30]

DURC

Adrian Costigan

25yd Prone Rifle

Place behind first bull of target

Round 1 Due Date : [July 27]

DURC

Ray Reilly

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]

DURC

Ray Reilly

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 2 Due Date : [August 10]

DURC

Ray Reilly

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 2 Due Date : [August 10]

DURC

Ray Reilly

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 3 Due Date : [August 24]

DURC

Ray Reilly

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 3 Due Date : [August 24]

DURC

Ray Reilly

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 4

Due Date : [September 7] DURC

Ray Reilly

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 4

Due Date : [September 7] DURC

Ray Reilly

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 5

Due Date : [September 21] DURC

Ray Reilly

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 5

Due Date : [September 21] DURC

Ray Reilly

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 6 Due Date : [October 5]

DURC

Ray Reilly

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 6 Due Date : [October 5]

DURC

Ray Reilly

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 7 Due Date : [October 19]

DURC

Ray Reilly

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 7 Due Date : [October 19]

DURC

Ray Reilly

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 8 Due Date : [November 2]

DURC

Ray Reilly

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 8 Due Date : [November 2]

DURC

Ray Reilly

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 9

Due Date : [November 16] DURC

Ray Reilly

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 9

Due Date : [November 16] DURC

Ray Reilly

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 10

Due Date : [November 30] DURC

Ray Reilly

25yd Prone Rifle Trinity Top Ten 2001

Round 10

Due Date : [November 30] DURC

Ray Reilly

25yd Prone Rifle

Place behind first bull of target

Round 1 Due Date : [July 27]

RRPC

Kealan Symes

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]

RRPC

Kealan Symes

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 2 Due Date : [August 10]

RRPC

Kealan Symes

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 2 Due Date : [August 10]

RRPC

Kealan Symes

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 3 Due Date : [August 24]

RRPC

Kealan Symes

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 3 Due Date : [August 24]

RRPC

Kealan Symes

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 4

Due Date : [September 7]

RRPC

Kealan Symes

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 4

Due Date : [September 7]

RRPC

Kealan Symes

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 5

Due Date : [September 21]

RRPC

Kealan Symes

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 5

Due Date : [September 21]

RRPC

Kealan Symes

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 6 Due Date : [October 5]

RRPC

Kealan Symes

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 6 Due Date : [October 5]

RRPC

Kealan Symes

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 7 Due Date : [October 19]

RRPC

Kealan Symes

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 7 Due Date : [October 19]

RRPC

Kealan Symes

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 8 Due Date : [November 2]

RRPC

Kealan Symes

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 8 Due Date : [November 2]

RRPC

Kealan Symes

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 9

Due Date : [November 16]

RRPC

Kealan Symes

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 9

Due Date : [November 16]

RRPC

Kealan Symes

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 10

Due Date : [November 30]

RRPC

Kealan Symes

25yd Prone Rifle Trinity Top Ten 2001

Round 10

Due Date : [November 30]

RRPC

Kealan Symes

25yd Prone Rifle

Place behind first bull of target

Round 1 Due Date : [July 27]

RRPC

Dave Cullen

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]

RRPC

Dave Cullen

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 2 Due Date : [August 10]

RRPC

Dave Cullen

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 2 Due Date : [August 10]

RRPC

Dave Cullen

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 3 Due Date : [August 24]

RRPC

Dave Cullen

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 3 Due Date : [August 24]

RRPC

Dave Cullen

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 4

Due Date : [September 7]

RRPC

Dave Cullen

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 4

Due Date : [September 7]

RRPC

Dave Cullen

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 5

Due Date : [September 21]

RRPC

Dave Cullen

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 5

Due Date : [September 21]

RRPC

Dave Cullen

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 6 Due Date : [October 5]

RRPC

Dave Cullen

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 6 Due Date : [October 5]

RRPC

Dave Cullen

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 7 Due Date : [October 19]

RRPC

Dave Cullen

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 7 Due Date : [October 19]

RRPC

Dave Cullen

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 8 Due Date : [November 2]

RRPC

Dave Cullen

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 8 Due Date : [November 2]

RRPC

Dave Cullen

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 9

Due Date : [November 16]

RRPC

Dave Cullen

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 9

Due Date : [November 16]

RRPC

Dave Cullen

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 10

Due Date : [November 30]

RRPC

Dave Cullen

25yd Prone Rifle Trinity Top Ten 2001

Round 10

Due Date : [November 30]

RRPC

Dave Cullen

25yd Prone Rifle

Place behind first bull of target

Round 1 Due Date : [July 27]

RRPC

Mick Delaney

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]

RRPC

Mick Delaney

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 2 Due Date : [August 10]

RRPC

Mick Delaney

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 2 Due Date : [August 10]

RRPC

Mick Delaney

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 3 Due Date : [August 24]

RRPC

Mick Delaney

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 3 Due Date : [August 24]

RRPC

Mick Delaney

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 4

Due Date : [September 7]

RRPC

Mick Delaney

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 4

Due Date : [September 7]

RRPC

Mick Delaney

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 5

Due Date : [September 21]

RRPC

Mick Delaney

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 5

Due Date : [September 21]

RRPC

Mick Delaney

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 6 Due Date : [October 5]

RRPC

Mick Delaney

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 6 Due Date : [October 5]

RRPC

Mick Delaney

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 7 Due Date : [October 19]

RRPC

Mick Delaney

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 7 Due Date : [October 19]

RRPC

Mick Delaney

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 8 Due Date : [November 2]

RRPC

Mick Delaney

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 8 Due Date : [November 2]

RRPC

Mick Delaney

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 9

Due Date : [November 16]

RRPC

Mick Delaney

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 9

Due Date : [November 16]

RRPC

Mick Delaney

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 10

Due Date : [November 30]

RRPC

Mick Delaney

25yd Prone Rifle Trinity Top Ten 2001

Round 10

Due Date : [November 30]

RRPC

Mick Delaney

25yd Prone Rifle

Place behind first bull of target

Round 1 Due Date : [July 27]

RRPC

Richard Farrell

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]

RRPC

Richard Farrell

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 2 Due Date : [August 10]

RRPC

Richard Farrell

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 2 Due Date : [August 10]

RRPC

Richard Farrell

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 3 Due Date : [August 24]

RRPC

Richard Farrell

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 3 Due Date : [August 24]

RRPC

Richard Farrell

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 4

Due Date : [September 7]

RRPC

Richard Farrell

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 4

Due Date : [September 7]

RRPC

Richard Farrell

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 5

Due Date : [September 21]

RRPC

Richard Farrell

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 5

Due Date : [September 21]

RRPC

Richard Farrell

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 6 Due Date : [October 5]

RRPC

Richard Farrell

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 6 Due Date : [October 5]

RRPC

Richard Farrell

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 7 Due Date : [October 19]

RRPC

Richard Farrell

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 7 Due Date : [October 19]

RRPC

Richard Farrell

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 8 Due Date : [November 2]

RRPC

Richard Farrell

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 8 Due Date : [November 2]

RRPC

Richard Farrell

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 9

Due Date : [November 16]

RRPC

Richard Farrell

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 9

Due Date : [November 16]

RRPC

Richard Farrell

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 10

Due Date : [November 30]

RRPC

Richard Farrell

25yd Prone Rifle Trinity Top Ten 2001

Round 10

Due Date : [November 30]

RRPC

Richard Farrell

25yd Prone Rifle

Place behind first bull of target

Round 1 Due Date : [July 27]

RRPC

Nick West

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]

RRPC

Nick West

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 2 Due Date : [August 10]

RRPC

Nick West

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 2 Due Date : [August 10]

RRPC

Nick West

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 3 Due Date : [August 24]

RRPC

Nick West

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 3 Due Date : [August 24]

RRPC

Nick West

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 4

Due Date : [September 7]

RRPC

Nick West

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 4

Due Date : [September 7]

RRPC

Nick West

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 5

Due Date : [September 21]

RRPC

Nick West

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 5

Due Date : [September 21]

RRPC

Nick West

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 6 Due Date : [October 5]

RRPC

Nick West

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 6 Due Date : [October 5]

RRPC

Nick West

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 7 Due Date : [October 19]

RRPC

Nick West

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 7 Due Date : [October 19]

RRPC

Nick West

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 8 Due Date : [November 2]

RRPC

Nick West

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 8 Due Date : [November 2]

RRPC

Nick West

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 9

Due Date : [November 16]

RRPC

Nick West

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 9

Due Date : [November 16]

RRPC

Nick West

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 10

Due Date : [November 30]

RRPC

Nick West

25yd Prone Rifle Trinity Top Ten 2001

Round 10

Due Date : [November 30]

RRPC

Nick West

25yd Prone Rifle

Place behind first bull of target

Round 1 Due Date : [July 27]

RRPC

J. Geogshesan

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]

RRPC

J. Geogshesan

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 2 Due Date : [August 10]

RRPC

J. Geogshesan

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 2 Due Date : [August 10]

RRPC

J. Geogshesan

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 3 Due Date : [August 24]

RRPC

J. Geogshesan

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 3 Due Date : [August 24]

RRPC

J. Geogshesan

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 4

Due Date : [September 7]

RRPC

J. Geogshesan

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 4

Due Date : [September 7]

RRPC

J. Geogshesan

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 5

Due Date : [September 21]

RRPC

J. Geogshesan

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 5

Due Date : [September 21]

RRPC

J. Geogshesan

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 6 Due Date : [October 5]

RRPC

J. Geogshesan

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 6 Due Date : [October 5]

RRPC

J. Geogshesan

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 7 Due Date : [October 19]

RRPC

J. Geogshesan

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 7 Due Date : [October 19]

RRPC

J. Geogshesan

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 8 Due Date : [November 2]

RRPC

J. Geogshesan

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 8 Due Date : [November 2]

RRPC

J. Geogshesan

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 9

Due Date : [November 16]

RRPC

J. Geogshesan

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 9

Due Date : [November 16]

RRPC

J. Geogshesan

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 10

Due Date : [November 30]

RRPC

J. Geogshesan

25yd Prone Rifle Trinity Top Ten 2001

Round 10

Due Date : [November 30]

RRPC

J. Geogshesan

25yd Prone Rifle

Place behind first bull of target

Round 1 Due Date : [July 27]

RRPC

Dave Trumper

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]

RRPC

Dave Trumper

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 2 Due Date : [August 10]

RRPC

Dave Trumper

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 2 Due Date : [August 10]

RRPC

Dave Trumper

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 3 Due Date : [August 24]

RRPC

Dave Trumper

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 3 Due Date : [August 24]

RRPC

Dave Trumper

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 4

Due Date : [September 7]

RRPC

Dave Trumper

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 4

Due Date : [September 7]

RRPC

Dave Trumper

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 5

Due Date : [September 21]

RRPC

Dave Trumper

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 5

Due Date : [September 21]

RRPC

Dave Trumper

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 6 Due Date : [October 5]

RRPC

Dave Trumper

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 6 Due Date : [October 5]

RRPC

Dave Trumper

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 7 Due Date : [October 19]

RRPC

Dave Trumper

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 7 Due Date : [October 19]

RRPC

Dave Trumper

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 8 Due Date : [November 2]

RRPC

Dave Trumper

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 8 Due Date : [November 2]

RRPC

Dave Trumper

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 9

Due Date : [November 16]

RRPC

Dave Trumper

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 9

Due Date : [November 16]

RRPC

Dave Trumper

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 10

Due Date : [November 30]

RRPC

Dave Trumper

25yd Prone Rifle Trinity Top Ten 2001

Round 10

Due Date : [November 30]

RRPC

Dave Trumper

25yd Prone Rifle

Place behind first bull of target

Round 1 Due Date : [July 27]

RRPC

Bobby Carolan

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]

RRPC

Bobby Carolan

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 2 Due Date : [August 10]

RRPC

Bobby Carolan

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 2 Due Date : [August 10]

RRPC

Bobby Carolan

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 3 Due Date : [August 24]

RRPC

Bobby Carolan

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 3 Due Date : [August 24]

RRPC

Bobby Carolan

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 4

Due Date : [September 7]

RRPC

Bobby Carolan

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 4

Due Date : [September 7]

RRPC

Bobby Carolan

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 5

Due Date : [September 21]

RRPC

Bobby Carolan

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 5

Due Date : [September 21]

RRPC

Bobby Carolan

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 6 Due Date : [October 5]

RRPC

Bobby Carolan

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 6 Due Date : [October 5]

RRPC

Bobby Carolan

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 7 Due Date : [October 19]

RRPC

Bobby Carolan

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 7 Due Date : [October 19]

RRPC

Bobby Carolan

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 8 Due Date : [November 2]

RRPC

Bobby Carolan

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 8 Due Date : [November 2]

RRPC

Bobby Carolan

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 9

Due Date : [November 16]

RRPC

Bobby Carolan

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 9

Due Date : [November 16]

RRPC

Bobby Carolan

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 10

Due Date : [November 30]

RRPC

Bobby Carolan

25yd Prone Rifle Trinity Top Ten 2001

Round 10

Due Date : [November 30]

RRPC

Bobby Carolan

25yd Prone Rifle

Place behind first bull of target

Round 1 Due Date : [July 27]

RRPC

J. Carroll

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]

RRPC

J. Carroll

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 2 Due Date : [August 10]

RRPC

J. Carroll

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 2 Due Date : [August 10]

RRPC

J. Carroll

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 3 Due Date : [August 24]

RRPC

J. Carroll

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 3 Due Date : [August 24]

RRPC

J. Carroll

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 4

Due Date : [September 7] RRPC

J. Carroll

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 4

Due Date : [September 7] RRPC

J. Carroll

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 5

Due Date : [September 21] RRPC

J. Carroll

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 5

Due Date : [September 21] RRPC

J. Carroll

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 6 Due Date : [October 5]

RRPC

J. Carroll

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 6 Due Date : [October 5]

RRPC

J. Carroll

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 7 Due Date : [October 19]

RRPC

J. Carroll

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 7 Due Date : [October 19]

RRPC

J. Carroll

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 8 Due Date : [November 2]

RRPC

J. Carroll

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 8 Due Date : [November 2]

RRPC

J. Carroll

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 9

Due Date : [November 16] RRPC

J. Carroll

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 9

Due Date : [November 16] RRPC

J. Carroll

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 10

Due Date : [November 30] RRPC

J. Carroll

25yd Prone Rifle Trinity Top Ten 2001

Round 10

Due Date : [November 30] RRPC

J. Carroll

25yd Prone Rifle

Place behind first bull of target

Round 1 Due Date : [July 27]

RRPC

I. Manley

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]

RRPC

I. Manley

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 2 Due Date : [August 10]

RRPC

I. Manley

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 2 Due Date : [August 10]

RRPC

I. Manley

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 3 Due Date : [August 24]

RRPC

I. Manley

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 3 Due Date : [August 24]

RRPC

I. Manley

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 4

Due Date : [September 7] RRPC

I. Manley

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 4

Due Date : [September 7] RRPC

I. Manley

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 5

Due Date : [September 21] RRPC

I. Manley

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 5

Due Date : [September 21] RRPC

I. Manley

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 6 Due Date : [October 5]

RRPC

I. Manley

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 6 Due Date : [October 5]

RRPC

I. Manley

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 7 Due Date : [October 19]

RRPC

I. Manley

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 7 Due Date : [October 19]

RRPC

I. Manley

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 8 Due Date : [November 2]

RRPC

I. Manley

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 8 Due Date : [November 2]

RRPC

I. Manley

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 9

Due Date : [November 16] RRPC

I. Manley

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 9

Due Date : [November 16] RRPC

I. Manley

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 10

Due Date : [November 30] RRPC

I. Manley

25yd Prone Rifle Trinity Top Ten 2001

Round 10

Due Date : [November 30] RRPC

I. Manley

25yd Prone Rifle

Place behind first bull of target

Round 1 Due Date : [July 27]

DRC

J.K.Walsh

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 1 Due Date : [July 27]

DRC

J.K.Walsh

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 2 Due Date : [August 10]

DRC

J.K.Walsh

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 2 Due Date : [August 10]

DRC

J.K.Walsh

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 3 Due Date : [August 24]

DRC

J.K.Walsh

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 3 Due Date : [August 24]

DRC

J.K.Walsh

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 4

Due Date : [September 7]

DRC

J.K.Walsh

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 4

Due Date : [September 7]

DRC

J.K.Walsh

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 5

Due Date : [September 21]

DRC

J.K.Walsh

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 5

Due Date : [September 21]

DRC

J.K.Walsh

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 6 Due Date : [October 5]

DRC

J.K.Walsh

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 6 Due Date : [October 5]

DRC

J.K.Walsh

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 7 Due Date : [October 19]

DRC

J.K.Walsh

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 7 Due Date : [October 19]

DRC

J.K.Walsh

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 8 Due Date : [November 2]

DRC

J.K.Walsh

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 8 Due Date : [November 2]

DRC

J.K.Walsh

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 9

Due Date : [November 16]

DRC

J.K.Walsh

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 9

Due Date : [November 16]

DRC

J.K.Walsh

25yd Prone Rifle

Place behind first bull of target

Trinity Top Ten 2001

Round 10

Due Date : [November 30]

DRC

J.K.Walsh

25yd Prone Rifle Trinity Top Ten 2001

Round 10

Due Date : [November 30]

DRC

J.K.Walsh

25yd Prone Rifle

Place behind first bull of target